A rectangle is 5 times as long as it is wide. The perimeter is 60 feet. Find the dimensions
Answer by TimothyLamb(4379)

L = 5w
2L + 2w = 60
---
2L + 2w = 60
2(5w) + 2w = 60
10w + 2w = 60
12w = 60
---
w = 5 ft
L = 25 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php